OpenStudy (anonymous):

c^2-6cd+9d^2-4 Factor this expression as a difference of squares...

OpenStudy (anonymous):

are you a genius?

OpenStudy (agent47):

c(c-6d)+(3d-2)(3d+2) ... hmm this way doesn't work

Parth (parthkohli):

Of course it does ^

Parth (parthkohli):

\( \color{Black}{\Rightarrow (a + b)(a - b) = a^2 - b^2 \Longrightarrow (3d - 2)(3d + 2) = (3d)^2 - (2)^2}\)

OpenStudy (agent47):

No, there's also the front part

Parth (parthkohli):

OOPSIE

Parth (parthkohli):

Well anyway let's start by doing \(c(c - d) + (9d^2 - 4)\)

OpenStudy (agent47):

c^2-4-6cd+9d^2 (c-2)(c+2)-6cd+9d^2

Parth (parthkohli):

Oh yup

OpenStudy (agent47):

ok if we pull out a 3d: (c-2)(c+2)+3d(3d-2c)

OpenStudy (agent47):

still not working.. ohhhh complete the square.. lol

OpenStudy (anonymous):

I really appreciate your help Agent47 :)

OpenStudy (agent47):

c^2-6cd+9d^2-4 DUHHHH LOL I got it

OpenStudy (anonymous):

I love you chloe! lol

OpenStudy (agent47):

factor the first part: (c-3d)^2-4

OpenStudy (agent47):

(c-3d)^2-2^2

Parth (parthkohli):

AWWWYEA

OpenStudy (anonymous):

so is that the answer? I love you too Austin.

OpenStudy (anonymous):

YAY! :)